# OTAD: An Optimal Transport-Induced Robust Model for Agnostic Adversarial Attack

### v1

@@ V-H Plain network-based OTAD exhibits reduced robustness
OTAD is designed based on DNNs with residual connections, as the forward propagation of a residual network approximates an optimal transport map. By leveraging the regularity of the optimal transport map, the DNN-induced discrete optimal transport map can be made continuous by solving the convex integration problem, resulting in a mapping with local Lipschitz continuity. What happens if we remove the residual connections? We can conjecture that the OTAD based on a plain network might exhibit reduced robustness. To investigate the necessity of residual connections for OTAD, we test the performance of ResNet and plain networks on relatively simple datasets, such as MNIST and Fashion MNIST. We compare the test accuracy, OTAD’s accuracy on clean samples, and robust accuracy under Square Attack (Table XIII).
TABLE XII: Standard Accuracy of KNN and OTAD on MNIST and CIFAR10
@@
TABLE XIII: Performance of ResNet and plain network-based OTADs on MNIST and Fashion MNIST against Square Attack with $500$ queries ($\epsilon=3$ for MNIST and $\epsilon=2$ for Fashion MNIST).
@@ V-I The limitations of OTAD for complex data
We generate $50000$ training and $10000$ test data of dimension 128. The data for each class $k$ are sampling from a Gaussian random vector $\mathcal{N}(\boldsymbol{\mu}_{k},\Sigma)$, where $\boldsymbol{\mu}_{k}\in\mathbb{R}^{d}$ is uniformly sampling from the unit sphere. We change the variance from $0.1$ to $0.6$ to control the difficulty of the classification task. We train a 3-block ResNet on each synthetic dataset and record test accuracy. The performance of OTAD is shown in Table XIV.
@@
TABLE XIV: Performance of OTAD on the synthetic data against BPDA + PGD ($\epsilon=1$). Std means the variance of various synthetic datasets.
@@ VI Conclusion and discussion
DNNs are fragile to adversarial attacks. To address this issue, we developed a novel two-step model OTAD to achieve accurate training data fitting while preserving the local Lipschitz property. First, we train a DNN to obtain a discrete optimal transport map from the data to its features. Taking advantage of the regularity property inherent in the optimal transport map, we employ convex integration to interpolate the map while ensuring the local Lipschitz property. The CIP can be solved through QCP or neural networks for fast computation. Our model suits popular architectures such as ResNets and Transformers. Experimental results demonstrate the superior performance of OTAD compared to other robust models on diverse types of datasets well-trained by a DNN with unchanged latent dimensionality. There are some directions to explore in the future:
@@ Acknowledgments
This work has been supported by the National Key Research and Development Program of China [No. 2019YFA0709501 to S.Z.], the National Natural Science Foundation of China [Nos. 12326614, 12126605], the R&D project of Pazhou Lab (Huangpu) [No. 2023K0602 to S.Z.], and the CAS Project for Young Scientists in Basic Research [No. YSBR-034 to S.Z.].

### v2

@@ V-H Plain network-based OTAD exhibits reduced robustness
OTAD is designed based on DNNs with residual connections, as the forward propagation of a residual network approximates an optimal transport map. By leveraging the regularity of the optimal transport map, the DNN-induced discrete optimal transport map can be made continuous by solving the convex integration problem, resulting in a mapping with local Lipschitz continuity. What happens if we remove the residual connections? We can conjecture that the OTAD based on a plain network might exhibit reduced robustness. To investigate the necessity of residual connections for OTAD, we test the performance of ResNet and plain networks on relatively simple datasets, such as MNIST and Fashion MNIST. We compare the test accuracy, OTAD’s accuracy on clean samples, and robust accuracy under Square Attack (Table XVI).
TABLE XV: Standard Accuracy of KNN and OTAD on MNIST and CIFAR10
@@
TABLE XVI: Performance of ResNet and plain network-based OTADs on MNIST and Fashion MNIST against Square Attack with $500$ queries ($\epsilon=3$ for MNIST and $\epsilon=2$ for Fashion MNIST).
@@ V-I The limitations of OTAD for complex data
We generate $50000$ training and $10000$ test data of dimension 128. The data for each class $k$ are sampling from a Gaussian random vector $\mathcal{N}(\boldsymbol{\mu}_{k},\Sigma)$, where $\boldsymbol{\mu}_{k}\in\mathbb{R}^{d}$ is uniformly sampling from the unit sphere. We change the variance from $0.1$ to $0.6$ to control the difficulty of the classification task. We train a 3-block ResNet on each synthetic dataset and record test accuracy. The performance of OTAD is shown in Table XVII.
@@
TABLE XVII: Performance of OTAD on the synthetic data against BPDA + PGD ($\epsilon=1$). Std means the variance of various synthetic datasets.
@@ VI Conclusion and discussion
DNNs are fragile to adversarial attacks. To address this issue, we developed a novel two-step model OTAD to achieve accurate training data fitting while preserving the local Lipschitz property. First, we train a DNN to obtain a discrete optimal transport map from the data to its features. Taking advantage of the regularity property inherent in the optimal transport map, we employ convex integration to interpolate the map while ensuring the local Lipschitz property. The convex integration problem can be solved through optimization solver or neural networks for fast computation. Our model suits popular architectures such as ResNets and Transformers. Experimental results demonstrate the superior performance of OTAD compared to other robust models on diverse types of datasets well-trained by a DNN with unchanged latent dimensionality. There are some directions to explore in the future:
@@ Acknowledgments
This work has been supported by the Strategic Priority Research Program of the Chinese Academy of Sciences [No. XDB0680101 to S.Z.], the CAS Project for Young Scientists in Basic Research [No. YSBR-034 to S.Z.], and the National Natural Science Foundation of China [Nos. 32341013, 12326614, 12126605].